I am having trouble understanding, "solving systems of equations by graphing."
Ex:
x+y=6
x-y=4
Then, I started to try and solve and find the ordered pair that worked for both equations. I only got this far, then I get confused:
x+y=6
-x -x
______
y=-1x+6

x-y=4
-x -x
_______
-y=-1+4

Answer by KnightOwlTutor(293)

x+y=6
x-y=4
Add each colum and you get 2x (+y-y=0)=6+4
2x=10
divide both sides by two and x=5
if x is 5 then y must equal 1
5+1=6
5-1=4